OpenStudy (anonymous):

Patrick and Brooklyn are making decisions about their bank accounts. Patrick wants to deposit $300 as a principle amount, with an interest of 3% compounded quarterly. Brooklyn wants to deposit $300 as the principle amount, with an interest of 5% compounded monthly. Explain which method results in more money after 2 years. Show all work.

OpenStudy (anonymous):

@Kidthatbro8 @Keigh2015 @longboardman15 @uybuyvf @Owlcoffee @insa @automaticloveletter @SyedMohammed98 @Kash_TheSmartGuy @mukushla @yashiii

OpenStudy (kropot72):

You need to use the following formula \[\large A=P(1+\frac{r}{n})^{nt}\] where P is the principal, A is the amount after t years, r is the interest rate expressed as a decimal, and n is the number of compounding periods in a year.

OpenStudy (kropot72):

So plugging in the values for Patrick's investment \[\large A _{P}=300(1+\frac{0.03}{4})^{4\times2}\] and for Brooklyn's investment \[\large A _{B}=300(1+\frac{0.05}{12})^{12\times2}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (kropot72):

You're welcome :)